OpenStudy (holly00d1248):

Solve for x 332X+9=1400X+25

OpenStudy (texaschic101):

we need to get all the x terms on one side and everything else on the other side. 332x + 9 = 1400x + 25 first subtract 332 from both sides of the equal sign. Can you do that ?

OpenStudy (texaschic101):

oops...subtract 332x from both sides

OpenStudy (holly00d1248):

1068

OpenStudy (holly00d1248):

Alright i got it

OpenStudy (texaschic101):

like this.. 332x + 9 = 1400x + 25 --subtract 332x from both sides 332x - 332x + 9 = 1400x - 332x + 25 -- combine like terms 9 = 1068x + 25 you understanding this so far ?

OpenStudy (texaschic101):

then subtract 25 from both sides 9 - 25 = 1068x + 25 - 25 -- combine terms - 16 = 1068x -- divide both sides by 1068 -16/1068 = x -3/267 = x (reduce) -1/89 = x